The work done by an external force to move a -7.30 μC charge from point A to point B is 1.70×10−3 J .

If the charge was started from rest and had 4.80×10−4 J of kinetic energy when it reached point B, what must be the potential difference between A and B?
Answer #1

here,

charge , Q = - 7.3 uC = - 7.3 * 10^-6 C

the work done , W = 1.7 * 10^-3 J

the final kinetic energy , KEf = 4.8 * 10^-4 J

let the potential difference be V

using Work-energy theorm

Q * V + We = KEf - KEi

- 7.3 * 10^-6 * V + 1.7 * 10^-3 = 4.8 * 10^-4

solving for V

V= 167.1 V

the potential difference between point A and B is 167.1 V